OpenStudy (anonymous):

Convert θ=3pi/2 to a rectangular equation please.

OpenStudy (anonymous):

you mean from "polar " form? you have a line i believe, through the origin. all you need is the slope

OpenStudy (amistre64):

whnever we are presented with a constant value, the tends to mean we get a line

OpenStudy (amistre64):

tan 3pi/2 tends to be the slope for this

OpenStudy (anonymous):

do you know how to find the slope fo that line?

OpenStudy (anonymous):

yeah, that will do it!

OpenStudy (amistre64):

since slope = y/x and tan = y/x they tend to be synonymous

OpenStudy (anonymous):

This was all the info given

OpenStudy (amistre64):

yes, it was

OpenStudy (amistre64):

you have all the information you need to construct it now. at which point are you having troubles with?

OpenStudy (anonymous):

oh I see now. Thanks!